OpenStudy (anonymous):

there are approximately 126 calories in a 2 ounce serving of lean hamburger and approximately 189 calories in a 3 ounce serving. write a linear equation for the number of calories in lean hammburger in terms of the size of the serving. Use your equation to estimate the number of calories in a 5 ounce serving of lean hamburger

OpenStudy (anonymous):

C = 63x

OpenStudy (anonymous):

when x = 5 ounces C = 315 calorie

OpenStudy (anonymous):

thank u! but how did u arrive to C=63x?

OpenStudy (anonymous):

well 2 ounces of hamburger contain 126 calories so if you divide 126 by 2 you get the number of calories in 1 ounce . Also 3 ounces contain 189 calories and 189 / 3 = 63 calories. So in order to count calories in any number of hamburgers we multiply that number ( which we call 'x') by 63. so calories = 63 * x or C = 63x

OpenStudy (anonymous):

algebra can be looked on as being a kind of shorthand in which we designate letters to unknown quantities - usuualy use x y or z for this purpose but any letter will do.

OpenStudy (anonymous):

wow. that makes so much more sense! thank!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

*thanks

OpenStudy (anonymous):

ok - the three roots of those equations sre correct - sometimes the book, teacher are wrong!